OpenStudy (anonymous):

Research scientists need a certain type of bacteria to conduct an experiment. There are 2,500 bacteria in a certain culture. The culture grows at a rate of 25% daily. The scientists needs at least 10,000 bacteria to conduct an experiment. What is the least number of days they need to wait for the bacteria culture to reach a quantity of 10,000? (Hint: use a guess-and-check method to determine the lowest number of days that satisfy the requirement.) A. 5 days B. 6 days C. 7 days D. 10 days

OpenStudy (imstuck):

I got 7 days. I did it the long way, but here it is: I started out with 2500 and at the end of day 1 we had the initial 2500 + 25% of 2500 = 2500+625=3125. Then at the end of the second day I had 3125 + (25%*3125)=3125+781.25=3906.25. Go on like this and at the end of 7 days you will have 11,920.92